# UniGaussian: Driving Scene Reconstruction from Multiple Camera Models via Unified Gaussian Representations

### v1

@@ 3.3 UniGaussian for Driving Scene Reconstruction
Although it is feasible to separately optimize different 3D scene representations for different camera models, this increases the difficulty for deployment in real-world simulators and prevents each camera model from learning complementary information. In this work, we proposes to learn a unified 3D Gaussian representation from multiple camera models for driving scene reconstruction. An overview of our framework is depicted in Fig. 2.
@@
With input images from different cameras, we represent a driving scene with composite 3D Gaussians111 In this work, “unified Gaussian representation” refers to learning a unified model for multiple sensors, while “composite Gaussians” refers to scene decomposition in the same model, so they are not contradictory. . Following [33, 34], we decompose the scene into background Gaussians for static scenes, dynamic Gaussians for moving objects and sky Gaussians for distant regions. Besides, we employ LiDAR point clouds to initialize 3D Gaussians by accumulating all LiDAR frames and projecting points onto images for color extraction. This provides a better representation of the scene geometry and enables the optional LiDAR simulation in our framework.
@@
To account for the distortion of fisheye images, we applied a series of affine transformations to further process 3D Gaussians for fisheye cameras as introduced in Sec. 3.2 while pinhole cameras skip these transformations. These processed 3D Gaussians are rendered with a unified tile-based rasterizer. Then, to resolve the exposure difference between pinhole and fisheye cameras, we applied camera-dependent scaling and biases factors on the rendering images as [33]. This models appearance difference between pinhole and fisheye cameras. Besides, we also render other modalities, including depth maps, semantic maps and normal maps. To render these maps, each 3D Gaussian is added with the corresponding 3D logits and then 2D maps are obtained via $\alpha$-blending of these 3D logits in the rasterizer. Although LiDAR simulation may not be directly achieved by explicit 3D Gaussians, we can obtain point clouds from the rendering depth maps by extracting points based on real-world LiDAR parameters and scans. Similarly, intensities of point clouds can be obtained by generating intensity maps from the rasterizer and extracting the corresponding points. Note that, LiDAR simulation is optional and not the focus of this work though we can derive it from depth and intensity maps.
@@
Driving scene reconstruction usually requires millions of 3D Gaussians. To enhance the adaptive density control of these Gaussians, we employ the Markov Chain Monte Carlo (MCMC) sampling and relocation strategy following [10]. In specific, we consider 3D Gaussian densification and pruning as a deterministic state transition of MCMC samples and employ a relocation strategy to dynamically adjust the position of Gaussians while preserving sample probability. With the opacity and scaling regularization, this adaptive control encourages the removal of redundant Gaussians and produces more compact scene Gaussians. Please refer to [10] for more details.
@@
4 Experiments
Our UniGaussian relies on the proposed differentiable rendering method of 3DGS for fisheye cameras. Thus, in our experiments, we firstly verify the effectiveness of the proposed rendering method in Sec. 4.1 and Sec. 4.2 and then examine the efficacy of the unified framework in Sec. 4.3. Please refer to the supplementary material for more details.
4.1 Fisheye Rendering Geometric Error Analysis
Since the proposed method is a “convert-project” method, the conversion part contains some approximations. After the light rays are twisted by fisheye cameras, the distorted 3D Gaussian no longer follows an exact normal distribution. Hence, the purpose of this experiment is to analyze the rendering geometric error introduced by the approximation.
Flowchart for Analysis.
We employ the method depicted in Fig. 4 to analyze the rendering geometric error introduced by these approximations. First, a 3D Gaussian model is trained by using the images of a pinhole camera. Then, images of a fisheye camera are rendered in two different ways: the “convert-project” method as introduced above and the “project-convert” method that renders a pinhole image and distorts it to a fisheye image. The “project-convert” method does not have any approximation in the conversion part, so its rendering result is considered as the reference image. We use the “BICYCLE” and “GARDEN” of the Mip-NeRF360 dataset [1] in this experiment because these data have enough multi-view images and accurate camera poses that enable rendering from any point of view.
@@ 4.1 Fisheye Rendering Geometric Error Analysis
Results Discussion.
The results are presented in Table 1. From these results, we can see that the stretching of 3D Gaussians is effective for improving the quality of rendering. This proves the correctness of our rendering algorithm’s derivation as the approximations of our method does not degrade the image quality. Moreover, in our method, most of the rendering time is spent on affine transformations of 3D Gaussians, while rasterization only takes around 3 ms. Note that, the time consumption shown in Table 1 is for images with large resolutions, but for autonomous driving scenes, 3D Gaussians are projected within a frustum and the image resolution is less than $4358\times 2824$ so the rendering can be done in real time. Besides, although the 2nd order approximation of polar stretching ratio gives improvement on rendering quality, it greatly increases the rendering time.
@@
4.2 Driving Scene Fisheye Camera Simulation
In this experiment, we further verify the effectiveness of the proposed rendering method for fisheye camera simulation in driving scene novel view synthesis.
Dataset.
We conduct experiments on the KITTI-360 dataset [14]. KITTI-360 contains images captured from fisheye cameras in real-world driving scenes and provides camera calibration and ego-vehicle pose. We select a segment of 221 images (frame ids 227-447 from “2013_05_28_drive_0000_sync”) and evenly select the every 8th image as the test set while the others are used as the training set. In addition to comparing the global image quality, three local zones (A, B, C) are compared. As it is shown in Fig. 5, zone “A” (Blue) is the bottom area where objects appear usually very close to the camera, zone “B” (Green) is the sky with little texture where the geometry is not very accurate, and zone “C” (Red) is far away from the camera with the large distortion in this region.
@@ 4.2 Driving Scene Fisheye Camera Simulation
We present the quantitative results in Table 2 and the qualitative results in Fig. 6. Overall, our approach achieves the best rendering quality both globally and locally (at zones A, B and C), while maintaining a real-time rendering speed of 39 FPS. In comparison, 3DGS+Undistort performs poorly on fisheye simulation, which indicates that direct distortion is not a suitable choice. Besides, from Table 2 and Fig. 7, we can see that our method without stretching yields slightly worse image quality, especially for texture details, but achieves significantly faster rendering speed of 138 FPS. Also, skipping stretching increases the number of Gaussians by $40\%$, which increases the risk of GPU memory overflow. The reason is that small Gaussians are insensitive to changes in the distribution caused by light rays distortion and the optimizer tends to approximate the scene with a large number of small 3D Gaussians in exchange for better rendering quality.
@@
4.3 Multiple Camera Model Simulation
In this experiment, we evaluate our UniGaussian framework for driving scene simulation from multiple camera models. Note that, to the best of our knowledge, no existing driving scene reconstruction method simulates both pinhole and fisheye cameras in a unified framework. Thus, the compared state-of-the-art methods are used for the reference purpose only, while modifying these methods for fisheye camera simulation beyond the scope of this work.
Dataset.
We conduct our experiments on KITTI-360 [14], because it is a real-world autonomous driving dataset providing both pinhole and fisheye images, while the other commonly used datasets provide only one type of images. For evaluation, we select four sequences as [24]. Each sequence contains 64 frames and we select the every 4th frame as the test set while the others are used as the training set.
@@ 4.3 Multiple Camera Model Simulation
Results Analysis.
We present the quantitative results in Table 3 and visualize some rendering results in Fig. 8. As shown in Table 3, for pinhole camera simulation, for fair comparison with AlignMiF and UniSim-SF, we also report the results of our approach with pinhole camera only. We can see that both ours and ours(pinhole) achieve better results than the state-of-the-art driving scene reconstruction methods, such as HUGS, AlignMiF and UniSim-SF. As for fisheye camera simulation, we add our fisheye rendering method to HUGS (named HUGS++) because HUGS does not support fisheye cameras. The results in Table 3 show that our approach outperforms the modified method. Besides, as shown in Fig. 8, our approach generates better rendering images with more fine-grained details. For example, our approach renders the license plate, the fence, the taillight, etc., while HUGS++ generates worse rendering images. Moreover, we also visualize some semantic and depth maps in Fig. 8. We can see that our multimodality outputs achieve slightly better holistic 3D scene understanding.
@@
Ablation study.
The ablation studies of the proposed rendering method’s stretching and approximation operations are discussed in Sec. 4.1 and Sec. 4.2. In this experiment, we further evaluate other components of our approach. From Table 4, we can see that our approach with all components perform the best and the supervisions from different modalities, including depth, semantic and normal, are helpful for model optimization. These modalities also facilitate better holistic scene understanding as shown in Fig. 8. Besides, from the last row of Table 4, we can see that without the adaptive control and $\mathcal{L}_{reg}$, the PSNRs are slightly worse, which indicates the importance of learning a compact representation in driving scene reconstruction.
@@ 5 Conclusion
In this work, we propose a new framework for learning unified 3D Gaussians of driving scenes from multiple camera models and present a new differentiable rendering method of 3D Gaussians tailored for fisheye cameras. Our method enables holistic driving scene understanding by modeling multiple sensors and modalities. Our experimental results on real-worl autonomous driving dataset verify the effectiveness of the proposed method.
Limitation. Due to the nature of explicit 3D Gaussian representations, 3DGS may not provide sufficient details when the viewing distance is close to the observation areas. This issue is exacerbated when adapting 3DGS to fisheye cameras in driving scene reconstruction due to the capability of fisheye cameras to observe nearby vehicles and buildings along the road. Our future work aims to solve this limitation and develop a real-world autonomous driving simulator.
@@
This supplementary material is organized as follows: In Sec. 6, we present more details of the experimental setups; in Sec. 7, we provide the implementation details of our approach; in Sec. 8, we present results of novel view synthesis with lane shift; in Sec. 9, we discuss the details of our approach to Lidar simulation.
@@ 6.2 Experimental Setup for Driving Scene Fisheye Camera Simulation
The MEI camera model is the only model provided for the fisheye camera on KITTI-360, so to evaluate our method on both the MEI and Kannala-Brandt models, we convert the MEI model to the Kannala-Brandt model. In specific, the conversion is approximating $\theta_{d}(\theta)$ determined by Eq. (3) of the main paper with the series expansion Eq. (2) of the main paper. Note that, when $\theta=0$, the first derivative of Eq. (2) equals $1$, while the first derivative of Eq. (3) is $1/(1+\xi)$. To eliminate this difference, Eq. (3) is modified to:
$\small\theta_{d}=\arctan r_{d}=\arctan\left(\left(\chi+k_{1}\chi^{3}+k_{2}\chi^{5}\right)(1+\xi)\right).$ (21)
@@ 6.3 Experimental Setup for Multiple Camera Model Simulation
HUGS [33] is a state-of-the-art 3DGS-based method for driving scene reconstruction and 3D scene understanding. We therefore use it as a baseline method for comparison. We modify HUGS based on the official code and the paper and additionally use our differentiable rendering method to train HUGS for fisheye camera simulation. For the other compared driving scene simulation methods, we borrow the results from [24]. Note that, although there are some other driving scene simulation methods, they are mostly designed for pinhole camera simulation so we choose the representative HUGS [33], AlignMiF[24] and UniSim-SF[31, 24] for comparison in our experiments. Modifying existing state-of-the-art methods for fisheye camera simulation in autonomous driving beyond the scope of this work.
@@ 7 Implementation Details
We implement our approach using python and pytorch. Following [9, 33], we set the initial position learning rate to $1.6\times 10^{-4}$ with decay to $1.6\times 10^{-6}$ at the last iteration. Both the scaling and rotation learning rates are set to $0.001$, while the learning rate of the spherical harmonics feature is set to $0.0025$ and the opacity learning rate is $0.05$. Adaptive Density Control is performed every 100 iterations starting from 500 iterations. For the composite scene Gaussians, the background Gaussians are initialize with LiDAR point clouds, the dynamic Gaussians are randomly initialized, and the sky Gaussians are uniformly initialized in a distant (more than 100 meters) spherical region for simplicity but more advanced sky cubemaps [30] can be used. The training loss $\mathcal{L}$ of our approach is defined in Eq. (20) of the main paper. The details of these losses and regularization terms are listed as below.
@@
It is the depth loss computed between the rendering depth and the monocular depth $\mathcal{L}_{d-mono}$ or LiDAR depth $\mathcal{L}_{d-lidar}$. Although the depth derived from LiDAR is accurate, it can only supervise the masked region with the projected point clouds. On the other hand, the monocular depth is coarse, but it can supervise the whole depth map. Thus, we define $\mathcal{L}_{d}$ as:
@@
8 Novel View Synthesis with Lane Shift
In driving scene reconstruction, it is important to synthesize novel views after the lane shift (left or right) of the ego vehicle for real-world simulators. In this experiment, we further present the results of novel view synthesis with lane shift @ 3 meters of the ego vehicle on KITTI-360. From Table 5 and Fig. 9, we can see that our approach is able to synthesize high-quality rendered images for novel view synthesis with lane shift while HUGS++ generates significantly worse results. For example, the FID of our approach is significantly better than HUGS++ and the buildings and vehicles in the rendered images of our approach are much clearer and more complete.

### v2

@@ 3.3 UniGaussian for Driving Scene Reconstruction
Fig. 2 depicts an overview of our approach to learning unified Gaussian representations from multiple camera models for driving scene reconstruction.
@@
With input images from different cameras, we represent a driving scene with composite 3D Gaussians111 In this work, “unified Gaussian representation” refers to learning a unified model for multiple sensors, while “composite Gaussians” refers to scene decomposition in the same model. . Following [35, 36, 31], we decompose the scene into background Gaussians for static scenes, dynamic Gaussians for moving objects and sky Gaussians for distant regions. Besides, we employ LiDAR point clouds to initialize 3D Gaussians by accumulating all LiDAR frames and projecting points onto images for color extraction. This provides a better representation of the scene geometry and enables the optional LiDAR simulation in our framework.
@@
To account for the distortion of fisheye images, we applied a series of affine transformations to further process 3D Gaussians for fisheye cameras as introduced in Sec. 3.2 while pinhole cameras skip these transformations. These processed 3D Gaussians are rendered with a unified tile-based rasterizer. Then, to resolve the exposure difference between pinhole and fisheye cameras, we applied camera-dependent scaling and biases factors on the rendering images as [35]. This models appearance difference between pinhole and fisheye cameras. Besides, we also render other modalities, including depth maps, semantic maps, and normal maps. To render these maps, each 3D Gaussian is added with the corresponding 3D logits, and then 2D maps are obtained via $\alpha$-blending of these 3D logits in the rasterizer. Although LiDAR simulation may not be directly achieved by explicit 3D Gaussians, we can obtain point clouds from the rendering depth maps by extracting points based on real-world LiDAR parameters and scans. Similarly, the intensities of point clouds can be obtained by generating intensity maps from the rasterizer and extracting the corresponding points. Please refer to the supplementary material for more details of the optional LiDAR simulation.
@@
Driving scene reconstruction usually requires millions of 3D Gaussians. To enhance the adaptive density control of these Gaussians, we employ the Markov Chain Monte Carlo (MCMC) sampling and relocation strategy following [11]. Specifically, we consider 3D Gaussian densification and pruning as a deterministic state transition of MCMC samples and employ a relocation strategy to dynamically adjust the position of Gaussians while preserving sample probability. With the opacity and scaling regularization, this adaptive control encourages the removal of redundant Gaussians and produces more compact scene Gaussians. Please refer to [11] for more details.
@@
4 Experiments
In our experiments, we first verify the effectiveness of the proposed rendering method in Sec. 4.1 and Sec. 4.2 and then examine the efficacy of the unified framework in Sec. 4.3.
4.1 Fisheye Rendering Geometric Error Analysis
Relative Error Analysis.
Since the proposed method is a “convert-project” method, the conversion part contains some approximations. After the light rays are twisted by fisheye cameras, the distorted 3D Gaussian no longer follows an exact normal distribution. The tangential stretching ration $k_{\phi}$ does not contain any approximation, while $k_{\theta}$ includes the truncation error. When using the first-order approximation, the truncation error of $k_{\theta}$ is defined as:
$\epsilon_{k_{\theta}}(\theta,\Delta\theta)=\frac{1}{2!}\frac{d^{2}\theta_{d}}{d\theta^{2}}\Delta\theta+\frac{1}{3!}\frac{d^{3}\theta_{d}}{d\theta^{3}}\Delta\theta^{2}+\ldots.$ (21)
On the image plane, this can cause scale errors in the radial direction from the center of the image, so the relative error is defined as $\epsilon_{k_{\theta}}(\theta,\Delta\theta)\cos{\theta_{d}}$. However, our design is not significantly affected by this approximation. To verify this, we conduct the following experiment.
Flowchart for Analysis.
We employ the method depicted in Fig. 4 to analyze the rendering geometric error introduced by these approximations. First, a 3D Gaussian model is trained using the images of a pinhole camera. Then, images of a fisheye camera are rendered in two different ways: the “convert-project” method as introduced above and the “project-convert” method that renders a pinhole image and distorts it to a fisheye image. The “project-convert” method does not have any approximation in the conversion part, so its rendering result is considered as the reference image. We use the “BICYCLE” and “GARDEN” of the Mip-NeRF360 dataset [1] in this experiment because these data have enough multi-view images and accurate camera poses that enable rendering from any point of view.
@@ 4.1 Fisheye Rendering Geometric Error Analysis
The results are presented in Tab. 1. From these results, we can see that the stretching of 3D Gaussians is effective for improving the quality of rendering. This proves the correctness of our rendering algorithm’s derivation because the approximations do not degrade the image quality. Moreover, in our method, most of the rendering time is spent on affine transformations of 3D Gaussians, while rasterization only takes around 3 ms. Note that, the time consumption shown in Tab. 1 is for images with large resolutions, but for autonomous driving scenes, 3D Gaussians are projected within a frustum and the image resolution is less than $4358\times 2824$ so the rendering can be done in real time. Besides, although the 2nd order approximation of the polar stretching ratio gives an improvement on rendering quality, it greatly increases the rendering time.
@@
4.2 Driving Scene Fisheye Camera Simulation
In this experiment, we verify the effectiveness of the proposed fisheye rendering method on the KITTI-360 dataset [15]. We compare the image quality of both global image and three representative local zones, namely, A, B, and C, as illustrated in Fig. 5. Details of the experimental setup can be found in the supplementary material.
We present the quantitative results in Tab. 2 and the qualitative results in Fig. 6. Overall, our approach achieves the best rendering quality both globally and locally (at zones A, B and C), while maintaining real-time rendering speed. In comparison, 3DGS+Undistort performs poorly on fisheye simulation, which indicates that direct distortion is not a suitable choice. Besides, from Tab. 2 and Fig. 7, we can see that our method without stretching yields slightly worse image quality, especially for texture details, but achieves significantly faster rendering speed of 138 FPS. Also, skipping stretching increases the number of Gaussians by $40\%$, which increases the risk of GPU memory overflow. The reason is that small Gaussians are insensitive to changes in the distribution caused by light rays distortion and the optimizer tends to approximate the scene with a large number of small 3D Gaussians in exchange for better rendering quality.
@@ 4.2 Driving Scene Fisheye Camera Simulation
4.3 Multiple Camera Model Simulation
In this experiment, we evaluate our UniGaussian framework for driving scene simulation from multiple camera models. We conduct our experiments on KITTI-360 [15], because it is a real-world autonomous driving dataset that provides both pinhole and fisheye images, while the other commonly used datasets provide only one type of image.
As shown in Tab. 3, for pinhole camera simulation, for fair comparison with AlignMiF and UniSim-SF, we also report the results of our approach with pinhole camera only. We can see that both ours and ours(pinhole) achieve better results than the state-of-the-art driving scene reconstruction methods, such as HUGS, AlignMiF and UniSim-SF. As for fisheye camera simulation, we add our fisheye rendering method to HUGS (named HUGS++) because HUGS does not support fisheye cameras. The results in Tab. 3 show that our approach outperforms the modified method. Besides, as shown in Fig. 8, our approach generates better rendering images with more fine-grained details. For example, our approach renders the license plate, the fence, the taillight, etc., while HUGS++ generates worse rendering images. Moreover, we also visualize some semantic and depth maps in Fig. 8. We can see that our multimodality outputs achieve slightly better holistic 3D scene understanding.
@@ 4.3 Multiple Camera Model Simulation
4.4 Ablation study
The ablation studies of the stretching and approximation operations of the proposed rendering method are discussed in Sec. 4.1 and Sec. 4.2. In this experiment, we further evaluate other components of our approach. From Tab. 4, we can see that our approach with all components performs the best and the supervisions from different modalities, including depth, semantic, and normal, are helpful for model optimization. These modalities also facilitate better holistic understanding of driving scenes as shown in Fig. 8. Besides, from the last row of Tab. 4, we can see that without the adaptive control and $\mathcal{L}_{reg}$, the PSNRs are slightly worse, indicating the importance of learning a compact representation in driving scene reconstruction.
@@ 5 Conclusion
In this work, we present a new differentiable rendering method of 3D Gaussians tailored for fisheye cameras and propose a new framework for learning unified 3D Gaussians of driving scenes from multiple camera models. Our method enables holistic driving scene understanding by modeling multiple sensors and modalities. Our experimental results on real-world autonomous driving dataset verify the effectiveness of the proposed method.
@@
This supplementary material is organized as follows: In Sec. 6, we present more details of the experimental setups; in Sec. 7, we provide the implementation details of our approach; in Sec. 8, we present the experimental comparison with Fisheye-GS [16]; in Sec. 9, we present results of novel view synthesis with lane shift; in Sec. 10, we discuss the details of our approach to Lidar simulation; in Sec. 11, we discuss the limitation and future work of our approach.
@@ 6.2 Experimental Setup for Driving Scene Fisheye Camera Simulation
Dataset.
We conduct experiments on the KITTI-360 dataset [15]. KITTI-360 contains images captured from fisheye cameras in real-world driving scenes and provides camera calibration and ego-vehicle pose. We select a segment of 221 images (frame ids 227-447 from “2013_05_28_drive_0000_sync”) and evenly select every 8th image as the test set while the others are used as the training set. In addition to comparing the global image quality, three local zones (A, B, C) are compared. As shown in Fig. 5, zone “A” (Blue) is the bottom area where objects appear usually very close to the camera, zone “B” (Green) is the sky with little texture where the geometry is not very accurate, and zone “C” (Red) is far away from the camera with the large distortion in this region.
@@
The MEI camera model is the only model provided for the fisheye camera on KITTI-360, so to evaluate our method on both the MEI and Kannala-Brandt models, we convert the MEI model to the Kannala-Brandt model. Specifically, the conversion is an approximation of $\theta_{d}(\theta)$ determined by Eq. (3) of the main paper with the series expansion Eq. (2) of the main paper. Note that, when $\theta=0$, the first derivative of Eq. (2) equals $1$, while the first derivative of Eq. (3) is $1/(1+\xi)$. To eliminate this difference, Eq. (3) is modified as:
$\theta_{d}=\arctan r_{d}=\arctan\left(\left(\chi+k_{1}\chi^{3}+k_{2}\chi^{5}\right)(1+\xi)\right).$ (22)
@@ 6.3 Experimental Setup for Multiple Camera Model Simulation
Dataset.
We conduct our experiments on KITTI-360 [15], because it is a real-world autonomous driving dataset that provides both pinhole and fisheye images, while the other commonly used datasets provide only one type of image. For evaluation, we select four sequences as [25]. Each sequence contains 64 frames, and we select every 4th frame as the test set while the others are used as the training set. Note that, although there are fisheye datasets for autonomous driving, such as Woodscape [33], they are constructed for perception tasks (e.g., object detection, segmentation, etc.) rather than driving scene reconstruction, so they often only have discrete frames and lack continuous video sequence data. In contrast, KITTI-360 provides a complete real-world autonomous driving sensor suite, including fisheye, pinhole and LiDAR, so evaluation on KITTI-360 for multiple camera model simulation is comprehensive to verify the effectiveness of our approach.
Methods Alameda Berlin London Nyc Average
Ours 26.0 23.8 27.9 22.3 25.0
Fisheye-GS [16] 24.1 24.2 25.1 20.3 23.4
[…]
Compared Methods.
To the best of our knowledge, no existing driving scene reconstruction method simulates both pinhole and fisheye cameras in a unified framework. HUGS [35] is a state-of-the-art 3DGS-based method for driving scene reconstruction and 3D scene understanding. We therefore use it as a baseline method for comparison. We modify HUGS based on the official code and the paper and additionally use our differentiable rendering method to train HUGS for fisheye camera simulation. For the other compared driving scene simulation methods, we borrow the results from [25]. Note that, although there are some other driving scene simulation methods, they are mostly designed for pinhole camera simulation, so we choose the representative HUGS [35], AlignMiF[25] and UniSim-SF[32, 25] for comparison in our experiments. Modifying existing state-of-the-art methods for fisheye camera simulation in autonomous driving is beyond the scope of this work.
@@ 7 Implementation Details
We implement our approach using python and pytorch. Following [10, 35], we set the initial position learning rate to $1.6\times 10^{-4}$ with decay to $1.6\times 10^{-6}$ at the last iteration. Both the scaling and rotation learning rates are set to $0.001$, while the learning rate of the spherical harmonics feature is set to $0.0025$ and the opacity learning rate is set to $0.05$. Adaptive Density Control is performed every 100 iterations starting from 500 iterations. For the composite scene Gaussians, the background Gaussians are initialized with LiDAR point clouds, the dynamic Gaussians are randomly initialized, and the sky Gaussians are uniformly initialized in a distant (more than 100 meters) spherical region for simplicity, but more advanced sky cubemaps [31] can be used. The training loss $\mathcal{L}$ of our approach is defined in Eq. (20) of the main paper. The details of these losses and regularization terms are listed below.
@@
It is the depth loss computed between the rendering depth and the monocular depth $\mathcal{L}_{d-mono}$ or LiDAR depth $\mathcal{L}_{d-lidar}$. Although the depth derived from LiDAR is accurate, it can only supervise the masked region with the projected point clouds. On the other hand, the monocular depth is coarse but can supervise the whole depth map. Thus, we define $\mathcal{L}_{d}$ as:
@@
8 Experimental Comparison with Fisheye-GS
As discussed in Sec. 2.1, Fisheye-GS [16] is only based on ideal camera models with equidistant projection, which hinders its use in driving scene reconstruction because fisheye cameras in driving scene are usually generic models and have large FOVs. Since Fisheye-GS is a state-of-the-art method adapting 3DGS to fisheye cameras, it would be interesting to compare our differentiable fisheye rendering method with Fisheye-GS. To this end, we conduct experimental comparison on Zip-NeRF(fisheye) [2] with four large scenes, namely, Alameda, Berlin, London, and Nyc. These scenes are respectively captured with a fisheye lens of 180 degree. We select every 8th frame as the test set while the others are used as the training set. As shown in Tab. 5, we approach achieves better performance compared with Fisheye-GS. Specifically, in the Alameda, London and Nyc scenes, we approach achieves significantly better results, while in the Berlin scene, our approach achieves comparable performance compared with Fisheye-GS. On average, our approach yields PSNR of 25.0 dB while Fisheye-GS obtains PSNR of 23.4 dB. These results verify the superiority of our approach over Fisheye-GS.